# Use of quantality in nuclei and many-body systems

### v2

@@ Abstract
The use of quantality is discussed in the case of nuclei and other many-body systems such as atomic electrons. This dimensionless quantity is known to indicate when a many-body system behaves like a crystal or a quantum liquid. Its role is further analyzed, showing the emergence of a fundamental lengthscale, the limit radius, which corresponds to the hard-core of the nucleon-nucleon interaction in the case of nucleons, and to a value close to the Bohr radius in the case of atomic electrons. The occurrence of a cluster phase in nuclei is analyzed using the quantality through its relation to the localization parameter, allowing for the identification of both the number of nucleons and the density as control parameters for the occurrence of this phase. The relation of the quantality to the magnitude of the interaction also exhibits a third dimensionless parameter, monitoring the magnitude of the spin-orbit effect in finite systems, through the realization of the pseudo-spin symmetry. The impact of quantality on the spin-orbit effect is compared in various many-body systems. The role of quantality in the relative effect of the binding energy and the shell one is also analyzed in nuclei. Finally, additional dimensionless quantities are proposed from the generalization of the quantality. Nuclei are found to be exceptional systems because all their dimensionless quantities are close to the order of unity, at variance with other many-body systems.
@@ I Introduction
Understanding clusterization in nuclei as a different phase, compared to the QL one is relevant. The use of the localization parameter calculated both analytically with the Harmonic Oscillator (HO) approximation, and with covariant Energy Density Functionnals (EDF), allowed to understand nuclear clusterization as a nuclear transitional phase, compared to the usual nuclear QL one [11, 17]. A transition from a gas of particles to nuclear liquid was also related to the properties of the nucleon-nucleon interaction with the lattice calculations [18, 12]. More recently, Effective Field Theories (EFTs) also tackle the occurrence of cluster phase in nuclei [19]. Recent ab initio PGCM calculations also confirm clusteristation traces in light nuclei, especially in 16O [20].
@@
The present work aims to analyze the behavior of many-body systems with simple but powerful indicators, by progressively refining their quantum description. First, the role of the ZPE is discussed. Then, the inclusion of the dispersion of the wave-function of the constituents, from the solution of a e.g. Schrödinger equation is considered. As a third step, the relation with the spin-orbit effect is given, based on the non-relativistic reduction of the Dirac equation. It will be shown how quantality plays a pivotal role during all these three stages. In section II, the emergence of a fundamental lengthscale, the limit radius, is studied in relation to the ZPE and the quantality. The relation between the quantality and the localization parameter is discussed in section III, showing how the appearance of clusters in the low-density regime is recovered. Section IV relates the quantality to the dimensionless magnitude of the interaction, showing how the spin-orbit effect in finite systems can be deduced. Finally, a systematic study of possible dimensionless quantities is provided in Section V, generalizing the concept of quantality.
II The quantality, the zero point energy, and the limit radius
@@
It should be noted that in a more general manner, the quantality could be defined as the ratio of the mean values of the kinetic energy to the potential one:
@@
where Eq. (1) is the application to the phenomenological case with a given potential and kinetic energy value. Eq. (3) allows the calculation of the value of the quantality for any interaction and many-body method, by using the corresponding wave functions and operators. Such a control parameter might prove useful in Ab Initio approaches. For instace, it could be possible to monitor the degree of clusterisation as a function of the similarity renormalisation group evolution of the bare chiral potential [30, 31, 32], or of the many-body scheme at stake [33, 34, 20] . One can then study with these approaches if, e.g., the quantum liquid behavior of the system is in agreement with the calculated value. In the present work, we will keep the use of Eq. (1), as it already paves the way for numerous developments.
@@
Starting with the inter-constituent distance r0, it has a lower boundary rl, so that rrl : below this limit radius rl, the ZPE is large enough to overcome the potential energy and make the system unbound. The corresponding limit condition T0=V0 leads to
@@
Fig. 1 gives a view of the interplay of V0, r0, rl and the ZPE considering a schematic interconstituent interaction, in the three following cases: rrl, rrl and rrl, for which the system is unbound, at the boundary limit and bound, respectively.
@@
The equilibrium distance r0 remains larger than rl because 2. However, the smaller , the larger interconstituent distance r0, and the smaller the delocalization effect, because the constituents are further apart. Therefore, drives localization effect. However, it misses finite-size effects. This will be taken into account in section III.
In nuclei the limit radius, being of the order of the hard-core one, gives a straightforward reason why the hard-core is washed out: it corresponds to the minimal radius allowed by the ZPE, and in practice, the mean inter-nucleon distance is r 1.2 fm, which is significantly larger. Therefore, due to a delocalization effect monitored by the quantality, the saturation density is lower than the packing one, which would correspond to a nucleonic system with interconstituent distance rl, the size of the repulsive core of the interaction. As discussed by Mottelson, this is an important question to understand [37, 38]. More precisely, using the relation between the interconstituent distance and the density:
@@ III Quantality and the cluster phase in nuclei
where is the typical spatial extension of the constituent wavefunction and r0 the interconstituent distance at equilibrium. In order to study cluster phases, this dimensionless quantity is 1 for homogeneous quantum liquid phase, 1 in the crystal case and 1 in the transitional cluster case. This quantity is related to the so-called Brückner one in condensed matter (see Appendix A). The interest of the localization parameter is that it takes into account finite-size effects, compared to quantality, by introducing as an additional quantity to consider in addition to m, r0 and V0.
In the case of nuclei, the spatial extension of the nucleonic wave function can be calculated microscopically from e.g. EDF approaches [14]. It can also be calculated from ab-initio or lattice EFT approaches for instance, by computing the spatial dispersion r of the corresponding wave functions. Similarly, the interconstituent distance could be computed from lattice-EFT calculations, in order to obtain a microscopic value. Here, in order to analyse the main impacts of the localization parameter, we use the sound approximation by the confining length of a harmonic oscillator (HO) potential, which parameters are chosen to match the radius of a given nucleus [14]:
@@
As an in-medium monitor for cluster phase occurrence where finite-size effect can be important, such as in nuclei or for atomic electrons, the localization parameter (10) can be generalized to any interconstituent distance :

### v3

@@ Abstract
The use of quantality is discussed in the case of nuclei and other many-body systems such as atomic electrons. This dimensionless quantity is known to indicate when a many-body system behaves like a crystal or a quantum liquid. Its role is further analyzed by showing its relation to the scattering length. The emergence of a fundamental lengthscale, the limit radius, is also shown. It corresponds to the hard-core of the nucleon-nucleon interaction in the case of nucleons, and to a value close to the Bohr radius in the case of atomic electrons. The occurrence of a cluster phase in nuclei is analyzed using the quantality through its relation to the localization parameter, allowing for the identification of both the number of nucleons and the density as control parameters for the occurrence of this phase. The relation of the quantality to the magnitude of the interaction also exhibits a third dimensionless parameter, monitoring the magnitude of the spin-orbit effect in finite systems, through the realization of the pseudo-spin symmetry. The impact of quantality on the spin-orbit effect is compared in various many-body systems. The role of quantality in the relative effect of the binding energy and the shell one is also analyzed in nuclei. Finally, additional dimensionless quantities are proposed from the generalization of the quantality. Nuclei are found to be exceptional systems because all their dimensionless quantities are close to the order of unity, at variance with other many-body systems.
@@ I Introduction
Therefore, understanding clusterization in nuclei as a different phase, compared to the QL one is relevant. The use of the localization parameter calculated both analytically with the Harmonic Oscillator (HO) approximation, and with covariant Energy Density Functionnals (EDF), allowed to understand nuclear clusterization as a nuclear transitional phase, compared to the usual nuclear QL one [11, 17]. A transition from a gas of particles to nuclear liquid was also related to the properties of the nucleon-nucleon interaction with the lattice calculations [18, 12]. More recently, Effective Field Theories (EFTs) also tackle the occurrence of cluster phase in nuclei [37]. Recent Ab initio PGCM calculations also confirm clusteristation traces in light nuclei, especially in 16O [20].
@@
The present work aims to analyze the behavior of many-body systems with simple but powerful indicators, by progressively refining their quantum description. First, the role of the ZPE is discussed. Then, the inclusion of the dispersion of the wave-function of the constituents, from the solution of a e.g. Schrödinger equation is considered. As a third step, the relation with the spin-orbit effect is given, based on the non-relativistic reduction of the Dirac equation. It will be shown how quantality plays a pivotal role during all these three stages. In section II, the study of the occurence of bound states from a given interaction first provides a relation between the quantality and the scattering length. Then, the emergence of a fundamental lengthscale, the limit radius, is studied in relation to the ZPE and the quantality. The relation between the quantality and the localization parameter is discussed in section III, showing how the appearance of clusters in the low-density regime is recovered. Section IV relates the quantality to the dimensionless magnitude of the interaction, showing how the spin-orbit effect in finite systems can be deduced. Finally, a systematic study of possible dimensionless quantities is provided in Section V, generalizing the concept of quantality.
II Quantality and bound states
II.1 General definition of quantality
@@
In a more general manner, the quantality could be defined as the ratio of the mean values of the kinetic energy to the potential one:
@@
where Eq. (1) is the application to the phenomenological case with a given potential and kinetic energy values. Eq. (3) allows the calculation of the value of the quantality for any interaction and many-body method, by using the corresponding wave functions and operators. Such a control parameter might prove useful in Ab Initio approaches. For instance, it could be possible to monitor the degree of clusterisation as a function of the similarity renormalisation group evolution of the bare chiral potential [30, 31, 32], or of the many-body scheme at stake [33, 34, 20]. One can then study with these approaches if, e.g., the quantum liquid behavior of the system is in agreement with the calculated value. In the present work, we will keep the use of Eq. (1), as it already paves the way for numerous developments.
II.2 Quantality and the scattering length
A first way to analyze the occurence of bound states from a given interaction is to relate the quantality to the scattering length. Let us consider an inter-constituent distance r0 and describe the two-body interaction by a spherically symmetric finite well. One can then derive an analytical expression of the scattering length [35]. Using the present notations and the definition (1) of the quantality, the scattering length becomes:
The behavior of a/r0 as a function of the quantality is shown in Fig. 1. The unitarity limit, for which the scattering length diverges, is obtained for specific values of , namely from Eq. (4):
where n=0,1,2, … The first unitarity limit corresponds to 0.4 and the second one for 0.05. The former value of the quantality is in the QL regime, whereas the latter is in the crystal one (see Table 1). In the case of the nucleon-nucleon interaction, it is striking that the corresponding quantality value of 0.4 allows to recover a scattering length close to the unitarity limit, in agreement with the phenomenology (a -19 fm in the case of the neutron-neutron interaction in the s-wave [36]). It should be noted that recent Ab initio calculations analyzed how clusterization emerges at the unitarity limit [37].
II.3 Quantality and the limit radius
@@
The inter-constituent distance r0, has also a lower boundary rl: below this limit radius rl, the ZPE is large enough to overcome the potential energy and make the system unbound. The corresponding limit condition T0=V0 leads to
@@
Fig. 2 gives a view of the interplay of V0, r0, rl and the ZPE considering the spherically symmetric finite well approximation of the inter-constituent interaction, in the three following cases: rrl, rrl and rrl, for which the system is unbound, at the boundary limit and bound, respectively.
@@
The equilibrium distance r0 remains larger than rl because 2. However, the smaller , the larger inter-constituent distance r0, and the smaller the delocalization effect, because the constituents are further apart. Therefore, drives localization effect. However, it misses finite-size effects. This will be taken into account in section III.
In nuclei the limit radius, being of the order of the hard-core one, gives a straightforward reason why the hard-core is washed out: it corresponds to the minimal radius allowed by the ZPE, and in practice, the mean inter-nucleon distance is r 1.2 fm, which is significantly larger. Therefore, due to a delocalization effect monitored by the quantality, the saturation density is lower than the packing one, which would correspond to a nucleonic system with inter-constituent distance rl, the size of the repulsive core of the interaction. As discussed by Mottelson, this is an important question to understand [36, 40]. More precisely, using the relation between the inter-constituent distance and the density:
@@ III Quantality and the cluster phase in nuclei
where is the typical spatial extension of the constituent wavefunction and r0 the inter-constituent distance at equilibrium. In order to study cluster phases, this dimensionless quantity is 1 for homogeneous quantum liquid phase, 1 in the crystal case and 1 in the transitional cluster case. This quantity is related to the so-called Brückner one in condensed matter (see Appendix A). The interest of the localization parameter is that it takes into account finite-size effects, compared to quantality, by introducing as an additional quantity to consider in addition to m, r0 and V0.
In the case of nuclei, the spatial extension of the nucleonic wave function can be calculated microscopically from e.g. EDF approaches [14]. It can also be calculated from ab-initio or lattice EFT approaches for instance, by computing the spatial dispersion r of the corresponding wave functions. Similarly, the inter-constituent distance could be computed from lattice-EFT calculations, in order to obtain a microscopic value. Here, in order to analyze the main impacts of the localization parameter, we use the sound approximation by the confining length of a harmonic oscillator (HO) potential, which parameters are chosen to match the radius of a given nucleus [14]:
@@
As an in-medium monitor for cluster phase occurrence where finite-size effect can be important, such as in nuclei or for atomic electrons, the localization parameter (12) can be generalized to any inter-constituent distance :